A fossil containing radioactive material has 25% of the original carbon-14 present. The half-life of carbon-14 is 5730 years. How old is the fossil?
1
Expert's answer
2020-11-24T05:55:01-0500

Solution.

"t(1/2) = \frac{ln(2)}{k}"

"k = 1.21 \times 10^{-4} \ years^{-1}"

"k \times t = ln(\frac{100}{25})"

t = 11457 years

Answer:

t = 11457 years